# MetaMetrics: Calibrating Metrics For Generation Tasks Using Human Preferences

### v2

@@ 6 Analysis and Discussion
In this section, we analyze the advantages of the methods in terms of interpretability, efficiency, and robustness. Additionally, we discuss how MetaMetrics enhances the capability of model evaluation.
@@ 6.2 Efficiency
Table 4 highlights the compute efficiency and superior performance of MetaMetrics-RM. MetaMetrics-RM, utilizing GP, combines GRM-Gemma-2B, Internlm2-7B-Reward, Skywork-Reward-Llama-3.1-8B, Internlm2-1.8B-Reward, and GRM-Llama3-8B totaling approximately 27B parameters. Despite this relatively small compute memory footprint, it outperforms models more than 2$\times$ its size. This demonstrates that MetaMetrics-RM offers state-of-the-art accuracy using significantly more cost-effective models, making it an ideal choice for resource-constrained applications while delivering even better performance.
@@ 6.3 Robustness
We evaluate the robustness of MetaMetrics through cross-dataset experiments in image captioning, tuning on one dataset and testing on another. Despite variations in domain and content, our results consistently demonstrate the effectiveness of MetaMetrics. Figure 3 shows that MetaMetrics-Cap tuned on THumB 1.0 and tested on Flickr8k and vice-versa out performs all individual metrics. This robustness indicates that MetaMetrics is not overfitting to specific dataset characteristics, but rather learning to combine metrics in a way that aligns with general human judgments of caption quality. This generalization capability is crucial for real-world applications, where the evaluation metric may need to perform well on diverse and potentially out-of-domain data.
@@ 7 Related Work
Ensemble-based metrics combine multiple individual metrics to enhance evaluation robustness through averaging (Adiwardana et al., 2020) or heuristics (Phy et al., 2020). The Absolute-Rating Multi-Objective Reward Model (ArmoRM) (Wang et al., 2024b) is a SotA reward model designed for RLHF (Ouyang et al., 2022). ArmoRM uses absolute ratings across 19 objectives trained to predict human preference scores.
@@ 8 Conclusion
Limitations and Future Works
In our experiment, we utilize metrics that are commonly employed to evaluate each downstream task. However, due to limitations in computational resources, we exclude LLMs exceeding 10.7B parameters to ensure that our metrics can be executed on commercial GPU resources with a maximum memory capacity of 40GB. Additionally, we restrict our exploration of metrics to avoid exhaustively incorporating LLM-based metrics, given our capacity and resource constraints.
Looking ahead, there are significant opportunities for further exploration. We can investigate metrics across additional languages and in larger multilingual generation tasks, where we identify MetaMetrics for multilingual applications (Winata et al., 2019). Furthermore, exploring metrics in the context of speech modalities presents another promising avenue for research.
Appendix A Metrics
@@
A.1 Abstractive Text Summarization
@@
LLM-based Metrics: SummaQA (Scialom et al., 2019).
A.2 Machine Translation
@@ A.4 Image Captioning
A.5 Reward Model Scoring
@@
Table 5: Reward Models.
Appendix B Datasets
@@ B.1 Data Statistics
Table 6: Dataset statistics used in the experiments.
B.2 Data Preprocessing
@@ Appendix C Optimization
C.1 Method Details
C.1.1 Matern Kernel
@@
C.1.2 Iterative-based Pruning
@@
C.2 Hyper-parameters
C.2.1 Bayesian Optimization
@@
Table 7: Hyper-parameters used for Gaussian Process on each task.
C.2.2 Boosting Method
@@
Metric SummEval Benchmark LLM Combined
Coh. Cons. Fluency Rel. Faith. Coh. Cons. Coh. Cons. Avg.
BLEU 0.110 0.126 0.113 0.170 0.162 0.246 0.252 0.029 0.082 0.143
chrF 0.143 0.094 0.071 0.198 0.272 0.163 0.342 0.068 0.218 0.174
METEOR 0.077 0.102 0.072 0.162 0.251 0.182 0.331 0.033 0.164 0.153
ROUGE-WE1 0.115 0.088 0.081 0.169 0.223 0.218 0.292 0.198 0.272 0.184
ROUGE-1 0.112 0.103 0.061 0.189 0.195 0.252 0.290 0.198 0.248 0.183
ROUGE-3 0.099 0.115 0.048 0.154 0.138 0.142 0.225 0.112 0.191 0.136
ROUGE-4 0.096 0.122 0.067 0.143 0.138 0.085 0.200 0.043 0.148 0.116
ROUGE-l 0.110 0.087 0.069 0.157 0.137 0.270 0.239 0.217 0.210 0.166
BARTScore (Max) 0.084 0.065 0.048 0.135 0.080 0.216 0.166 0.054 0.057 0.101
BARTScore (Mean) 0.086 0.074 0.040 0.143 0.080 0.216 0.166 0.210 0.173 0.132
BLEURT (Max) 0.185 0.070 0.114 0.189 0.046 0.264 0.143 0.229 0.143 0.154
BLEURT (Mean) 0.099 0.011 0.029 0.130 0.046 0.264 0.143 0.265 0.173 0.129
BERTScore (f1) 0.105 0.100 0.120 0.181 0.171 0.280 0.283 0.306 0.304 0.206
BERTScore (Recall) 0.104 0.111 0.093 0.176 0.228 0.223 0.333 0.218 0.300 0.198
SummaQA (F1) 0.051 0.120 0.106 0.114 0.318 0.038 0.293 -0.014 0.180 0.134
SummaQA (Prob) 0.070 0.100 0.070 0.138 0.330 -0.003 0.303 -0.028 0.206 0.132
@@
GP (Full) 0.170 0.142 0.115 0.233 0.356 0.308 0.383 0.315 0.333 0.262
GP (Top 5) 0.149 0.137 0.117 0.211 0.342 0.307 0.370 0.296 0.316 0.249
GP (Oracle Top 5) 0.169 0.138 0.121 0.230 0.356 .308 0.384 0.296 0.316 0.258
XGBoost 0.190 0.151 0.131 0.271 0.406 0.411 0.397 0.403 0.367 0.303
XGBoost (Iterative Best) 0.185 0.165 0.181 0.255 0.406 0.404 0.394 0.402 0.369 0.307
XGBoost (Iterative Top 5) 0.141 0.115 0.160 0.208 0.377 0.378 0.383 0.369 0.351 0.276
XGBoost (Top 5) 0.110 0.096 0.120 0.186 0.314 0.353 0.369 0.343 0.327 0.246
XGBoost (Oracle Top 5) 0.135 0.090 0.134 0.219 0.319 0.354 0.378 0.343 0.325 0.255
Table 9: Test result for summarization task. Coh., Cons., Rel., and Faith. corresponds to coherence, consistency, relevance, and faithfulness respectively. Bold and underlined values indicate the best and second best performance, respectively. The Combined column reflects results when all datasets are merged.
@@
Appendix D Detailed Results
@@
D.1 RewardBench Leaderboard
The leaderboard ranking screenshot is shown in Figure 5.
Figure 5: RewardBench Leaderboard Screenshot (Updated on October 6th, 2024).
Figure 6: RewardBench Detailed Leaderboard Screenshot (Updated on October 6th, 2024).
@@ D.1 RewardBench Leaderboard
D.2 Abstractive Text Summarization
Table 9 provides a detailed breakdown of the correlations between various metrics and human preference measurements for abstractive text summarization.
D.3 Machine Translation
@@ D.5 Image Captioning
D.6 Reward Model Scoring
@@
GP 93.4 98.3 86.4 90.8 98.2
GP (Top 3) 93.0 95.5 87.7 91.6 97.0
XGBoost 92.8 96.1 88.4 90.9 95.8
XGBoost (Iterative Best) 92.9 95.8 88.4 91.8 95.8
XGBoost (Top 3) 93.0 86.0 90.3 93.0 95.5
Appendix E Weights and Feature Importance
E.1 Question Answering

### v3

@@ 6 Analysis and Discussion
In this section, we analyze the advantages of the methods in terms of interpretability, efficiency, and robustness. We discuss how MetaMetrics enhances the capability of model evaluation.
@@ 6.2 Effectiveness and Efficiency
Table 4 highlights the compute efficiency and superior performance of MetaMetrics-RM. MetaMetrics-RM, utilizing GP. Despite this relatively small compute memory footprint, it outperforms models with 70B parameters. This demonstrates that MetaMetrics-RM offers state-of-the-art accuracy using significantly more cost-effective models, making it an ideal choice for resource-constrained applications while delivering even better performance.
@@ 6.3 Robustness
We present our robustness evaluation in both cross-lingual and cross-dataset scenarios. Comprehensive experimental results are detailed in Subsection F.7. For the cross-lingual evaluation, we assess our method on WMT23 and WMT24 using unseen language pairs. Specifically, we evaluate MetaMetrics on he-en (Hebrew-English), en-es (English-Spanish), and ja-zh (Japanese-Chinese). Overall, our approach outperforms all existing baseline metrics. We also evaluate the robustness of MetaMetrics through cross-dataset experiments in image captioning, tuning on one dataset and testing on another. Despite variations in domain and content, our results consistently demonstrate the effectiveness of MetaMetrics. Figure 2 shows that MetaMetrics-Cap tuned on THumB 1.0 and tested on Flickr8k and vice-versa out performs all individual metrics. This robustness indicates that MetaMetrics is not overfitting to specific dataset characteristics, but rather learning to combine metrics in a way that aligns with general human judgments of caption quality. This generalization capability is crucial for real-world applications, where the evaluation metric may need to perform well on diverse and potentially out-of-domain data.
@@ 7 Related Work
Ensemble-based metrics combine multiple individual metrics to enhance evaluation robustness through averaging (Adiwardana et al., 2020) or heuristics (Phy et al., 2020). The Absolute-Rating Multi-Objective Reward Model (ArmoRM) (Wang et al., 2024b) is a SotA reward model designed for RLHF (Ouyang et al., 2022).
LLM-based Metrics.
LLM-based metrics leverage generative language models to assess the quality of an input by generating scores. Examples include BARTScore (Yuan et al., 2021), G-Eval (Liu et al., 2023) utilizing the GPT-4 model (Achiam et al., 2023), Unieval (Zhong et al., 2022) utilizing the T5 model (Raffel et al., 2020), MetricX (Juraska et al., 2023; 2024), COMET, CometKiwi (Rei et al., 2022), XCOMET (Guerreiro et al., 2023), and GEMBA-MQM (Kocmi & Federmann, 2023).
@@ 8 Conclusion
Acknowledgments
We would like to express our sincere gratitude to Mingyang Zhou for his insightful suggestions and discussions on this project.
Appendix A Limitations and Future Work
In our experiment, we utilize metrics that are commonly employed to evaluate each downstream task. However, due to limitations in computational resources, we exclude LLMs exceeding 10.7B parameters to ensure that our metrics can be executed on commercial GPU resources with a maximum memory capacity of 48GB. Additionally, we restrict our exploration of metrics to avoid exhaustively incorporating LLM-based metrics, given our capacity and resource constraints.
Looking ahead, there are significant opportunities for further exploration. We can explore into metrics across various promising research avenues, such as additional languages, expanded vision and speech modalities (Ho et al., 2020; Blattmann et al., 2023; Zheng et al., 2024), and larger multilingual generation tasks. In particular, we identify MetaMetrics as a valuable tool for multilingual (Winata et al., 2019a; Joshi et al., 2020; Anugraha et al., 2024), multi-cultural (Adilazuarda et al., 2024; Winata et al., 2024a; Yue et al., 2024), and code-switching (Winata et al., 2019b; 2023; Kuwanto et al., 2024) applications.
Appendix B Metrics
@@
B.1 Abstractive Text Summarization
@@
LLM-based Metrics: SummaQA (Scialom et al., 2019), UniEval (Zhong et al., 2022), G-Eval (Liu et al., 2023).
B.2 Machine Translation
@@ B.4 Image Captioning
B.5 Reward Model Scoring
@@
Appendix C Metric Design
C.1 Framework
Figure 4 illustrates the framework for both reference-free and reference-based settings.
C.2 Alternative Weighting Score
Multiplicative GP
Recall that in the linear-weighting scheme for GP, given a set of $N$ evaluation metrics $\hat{\mathbf{y}}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{N}]$ with their associated weights of $\{w_{1},w_{2},\ldots,w_{N}\}$, GP linear is given by $\hat{y}_{\textsc{MM-Linear}}(\mathbf{w})=\mathbf{w}^{\top}\hat{\mathbf{y}}$.
To explore more weighting schemes, we introduce “Multiplicative” GP, which accounts for interactions between metrics through pairwise products. For $N$ metrics, there are $c=\binom{N}{2}=\frac{N\cdot(N-1)}{2}$ unique pairwise combinations. Let $\mathbf{w_{\text{pair}}}$ be the weight vector for these combinations, defined as $\mathbf{w_{\text{pair}}}=[w_{1},w_{2},\ldots,w_{c}]$ and let $\hat{\mathbf{y}}_{\text{pair}}$ be the vector of pairwise products, where $\hat{\mathbf{y}}_{\text{pair}}=[\hat{y}_{i}\hat{y}_{j}\mid 1\leq i<j\leq N].$ The Multiplicative GP output is then computed as $\hat{y}_{\text{MM-multi}}(\mathbf{w_{\text{pair}}})=\mathbf{w_{\text{pair}}}^{\top}\hat{\mathbf{y}}_{\text{pair}}$.
Appendix D Datasets
@@ D.1 Data Statistics
D.2 Data Preprocessing
@@ Appendix E Optimization
E.1 Method Details
E.1.1 Matern Kernel
@@
E.1.2 Iterative-based Pruning
@@
E.2 Hyper-parameters
E.2.1 Bayesian Optimization
@@
Table 7: Hyper-parameters used for Gaussian Process on each task.
@@
E.2.2 Boosting Method
@@
Table 9: Kendall and spearman correlation detailed results with human ratings for summarization task on SummEval. Coh., Cons., Fluency, and Rel. corresponds to coherence, consistency, and fluency, relevance, respectively. Bold and underlined values indicate the best and second best performance, respectively.
Metric Kendall Spearman
Coh. Cons. Fluency Rel. Avg. Coh. Cons. Fluency Rel. Avg.
BLEU 0.110 0.126 0.113 0.170 0.130 0.157 0.160 0.145 0.239 0.175
chrF 0.143 0.094 0.071 0.198 0.127 0.205 0.119 0.091 0.278 0.173
METEOR 0.077 0.102 0.072 0.162 0.103 0.108 0.130 0.093 0.229 0.140
ROUGE-1 0.112 0.103 0.061 0.189 0.116 0.159 0.130 0.079 0.264 0.158
ROUGE-3 0.099 0.115 0.048 0.154 0.104 0.142 0.147 0.063 0.219 0.143
ROUGE-4 0.096 0.122 0.067 0.143 0.107 0.137 0.155 0.087 0.203 0.146
ROUGE-l 0.110 0.087 0.069 0.157 0.106 0.156 0.111 0.089 0.219 0.144
ROUGE-WE1 0.115 0.088 0.081 0.169 0.113 0.164 0.112 0.105 0.237 0.115
BLEURT (Max) 0.185 0.070 0.114 0.189 0.140 0.262 0.088 0.062 0.194 0.152
BLEURT (Mean) 0.099 0.011 0.029 0.130 0.067 0.123 0.094 0.052 0.202 0.118
BERTScore (f1) 0.105 0.100 0.120 0.181 0.127 0.150 0.128 0.155 0.256 0.172
BERTScore (Recall) 0.104 0.111 0.093 0.176 0.121 0.148 0.141 0.120 0.250 0.165
SummaQA (f1) 0.051 0.120 0.106 0.114 0.098 0.073 0.152 0.136 0.165 0.132
SummaQA (Prob) 0.070 0.100 0.070 0.138 0.095 0.098 0.128 0.092 0.196 0.129
LLM-based Metrics
BARTScore (Mean) 0.086 0.074 0.040 0.143 0.086 0.123 0.094 0.052 0.202 0.118
UniEval 0.413 0.353 0.359 0.324 0.362 0.577 0.439 0.458 0.446 0.480
G-Eval (GPT4) 0.429 0.413 0.409 0.437 0.422 0.565 0.510 0.470 0.581 0.531
Ensemble Baselines
Uniform 0.141 0.133 0.117 0.213 0.151 0.201 0.170 0.150 0.298 0.205
Weighted Avg 0.150 0.141 0.122 0.220 0.159 0.215 0.179 0.158 0.309 0.215
@@
   GP 0.172 0.140 0.130 0.252 0.174 0.244 0.179 0.167 0.354 0.236
   XGBoost 0.192 0.186 0.186 0.276 0.210 0.274 0.236 0.239 0.386 0.284
w/ LLM-based Metrics
   GP 0.454 0.419 0.409 0.449 0.433 0.609 0.519 0.470 0.601 0.550
   GP (Top 2) 0.461 0.428 0.409 0.449 0.437 0.628 0.528 0.470 0.601 0.557
   XGBoost 0.476 0.367 0.404 0.447 0.424 0.642 0.460 0.512 0.600 0.553
   XGBoost (Top 2) 0.476 0.436 0.430 0.445 0.447 0.636 0.508 0.511 0.594 0.562
Table 10: Kendall correlation results with human ratings for summarization task on Benchmark LLM. Faith., Coh., and Cons. correspond to faithfulness, coherence, and consistency, respectively. Bold and underlined values indicate the best and second best performance, respectively.
Metric Benchmark LLM
Faith. Coh. Cons.
BLEU 0.162 0.246 0.252
chrF 0.272 0.163 0.342
METEOR 0.251 0.182 0.331
ROUGE-1 0.195 0.252 0.290
ROUGE-3 0.138 0.142 0.225
ROUGE-4 0.138 0.085 0.200
ROUGE-l 0.137 0.270 0.239
ROUGE-WE1 0.223 0.218 0.292
BARTScore (Mean) 0.080 0.216 0.166
BLEURT (Max) 0.046 0.264 0.143
BLEURT (Mean) 0.046 0.264 0.143
BERTScore (f1) 0.171 0.280 0.283
BERTScore (Recall) 0.228 0.223 0.333
SummaQA (f1) 0.318 0.038 0.293
SummaQA (Prob) 0.330 -0.003 0.303
Ensemble Baselines
Uniform 0.226 0.234 0.328
Weighted Avg 0.256 0.264 0.337
MetaMetrics-Sum
GP 0.356 0.308 0.383
XGBoost 0.406 0.411 0.397
XGBoost (Iterative Top 5) 0.377 0.378 0.383
@@
Appendix F Detailed Results
@@
F.1 Abstractive Text Summarization
Table 9 provides a detailed breakdown of the correlations between various metrics and human preference measurements for abstractive text summarization on SummEval LLM. Table 10 provides a detailed breakdown of the correlations between various metrics and human preference measurements for abstractive text summarization on Benchmark LLM.
F.2 Machine Translation
@@ F.4 Image Captioning
F.5 Reward Model Scoring
@@
GP 93.5 98.9 86.2 90.7 98.2
GP (Top 3) 92.8 98.3 84.4 90.7 97.9
XGBoost 92.9 95.8 89.7 92.2 94.0
XGBoost (Iterative Best) 92.7 95.5 89.0 91.6 94.7
XGBoost (Iterative Top 3) 92.3 95.5 87.7 91.2 94.8
XGBoost (Top 3) 92.2 96.6 86.2 91.2 94.7
F.6 Ablation Study
Table 15: Ablation study on GP using linear, multiplicative, and combined (linear and multiplicative) weightings on Machine Translation (MT), and Reward Model tasks.
Method MT MT (QE) Reward Model
GP (Linear) 0.819 0.801 93.5
GP (Multiplicative) 0.822 0.798 91.2
GP (Combined) 0.818 0.803 92.1
Table 16: Kendall correlation results with human ratings for summarization task with Spearman and Kendall as the evaluation functions. Coh., Cons., Rel., and Faith. corresponds to coherence, consistency, relevance, and faithfulness respectively.
Metric SummEval Benchmark LLM Combined
Coh. Cons. Fluency Rel. Faith. Coh. Cons. Coh. Cons. Avg.
XGBoost (Kendall) 0.475 0.376 0.407 0.439 0.406 0.411 0.397 0.403 0.367 0.409
XGBoost (Spearman) 0.476 0.376 0.403 0.448 0.406 0.411 0.397 0.402 0.369 0.409
Table 15 presents the results of an ablation study on the application of GP using three different weighting methods: linear, multiplicative, and combined (linear and multiplicative) on Summarization, Machine Translation, and Reward Model tasks as described in C. The results indicate that the linear weighting method generally outperforms the other approaches on average. This is likely due to GP’s preference for operating on lower-dimensional spaces Li et al. (2016); Frazier (2018); Nayebi et al. (2019); Malu et al. (2021); Binois & Wycoff (2022).These findings reiterate the benefits of using fewer metrics for evaluation, which simplifies the optimization process.
Table 16 compares Kendall and Spearman correlations with human ratings for the summarization task using XGBoost. The results show minimal performance differences between the two objective functions, which is expected given the similarity of evaluation measurement between Spearman and Kendall correlations.
F.7 Robustness
We demonstrate the robustness of MetaMetrics in two scenarios: cross-lingual and cross-dataset. The cross-lingual setting assesses our method’s capability with unseen languages, while the cross-dataset setting evaluates its performance in the face of distribution shifts or when the dataset originates from a different domain.
F.7.1 Cross-lingual Setting
We evaluate our method on WMT23 and WMT24 for unseen language pairs. To calibrate MetaMetrics, we employ three specific language pairs: zh-en (Chinese-English), en-de (English-German), and en-ru (English-Russian). For WMT23, we test our calibrated metric on unseen language pairs, such as he-en (Hebrew-English), with results presented in Table 17. For WMT24, we extend the evaluation to include en-es (English-Spanish) and ja-zh (Japanese-Chinese), with results detailed in Table 18.
Table 17: Results on Cross-lingual settings on WMT23 for Hebrew-English (he-en). Bold and underline values indicate the best and second best performance, respectively.
Metric he-en
sys seg seg
pearson pearson acc-t
chrF 0.776 0.221 0.460
BLEU 0.769 0.220 0.442
BERTScore 0.895 0.335 0.515
Yisi-1 0.917 0.395 0.529
MetricX-23-XXL 0.910 0.548 0.577
XCOMET-Ensemble 0.950 0.556 0.586
COMET 0.940 0.401 0.531
XCOMET-XL 0.949 0.530 0.576
XCOMET-XXL 0.952 0.465 0.560
MetaMetrics-MT
   GP 0.947 0.546 0.590
   XGBoost 0.957 0.557 0.574
Table 18: Average Results of acc-t and Soft-pairwise accuracy (SPA) on Cross-lingual settings on WMT24 (Freitag et al., 2024) for English-Spanish (en-es) and Japanese-Chinese (ja-zh). We report the number following the metric calculated by WMT24 organizers for language breakdown performance since the test set labels are unseen to the authors. Bold and underline values indicate the best and second best performance, respectively.
Metric en-es ja-zh avg
sentinel-ref-mqm 0.631 0.490 0.561
BLEU 0.596 0.588 0.592
spBLEU 0.602 0.590 0.596
chrF 0.615 0.616 0.616
chrfS 0.630 0.602 0.616
BERTScore 0.594 0.651 0.623
MEE4 0.635 0.625 0.630
damonmonli 0.688 0.633 0.661
YiSi-1 0.657 0.666 0.662
PrismRefSmall 0.666 0.667 0.667
PrismRefMedium 0.734 0.545 0.640
BLCOM_1 0.698 0.676 0.687
BLEURT-20 0.688 0.685 0.687
COMET-22 0.744 0.636 0.690
XCOMET 0.740 0.700 0.720
MetricX-24 (Hybrid) 0.742 0.718 0.730
MetaMetrics-MT (GP) 0.745 0.717 0.731
F.7.2 Cross-Dataset Setting
In addition to the cross-lingual setting, we also explore a cross-dataset scenario specifically for image captioning, as shown in Table 19. Here, we present the results when MetaMetrics-Cap is tuned on THuMB 1.0 and tested on Flickr8k, and vice versa. Our findings reveal that MetaMetrics-Cap, whether using GP or XGBoost, outperforms individual metrics, demonstrating its robustness in the presence of data shifts.
Table 19: Kendall correlation Cross-Dataset results of MetaMetrics-Cap when it is tuned on THumB 1.0 and tested on Flickr8k, and vice versa. The columns correspond to the testing dataset. Scores are reported for multiple metrics, with the average performance (avg.) across Cross-Dataset results. Bold and underlined values indicate the best and second best performance, respectively.
Metric Flickr8k THumB 1.0 avg.
BLEU-1 0.377 0.264 0.321
BLEU-4 0.361 0.210 0.286
ROUGE 0.375 0.237 0.306
METEOR 0.481 0.281 0.381
CIDEr 0.498 0.263 0.381
SPICE-F 0.591 0.200 0.396
SPICE-PR 0.603 0.170 0.387
SPICE-RE 0.579 0.221 0.400
TIGEr 0.542 0.153 0.348
CLIP-S 0.514 0.292 0.403
RefCLIP-S 0.576 0.210 0.393
MetaMetrics-Cap
GP (All) 0.638 0.298 0.468
GP (Top 5) 0.560 0.246 0.403
XGBoost (All) 0.595 0.323 0.449
XGBoost (Top 5) 0.575 0.311 0.443
XGBoost (Iterative-Best) 0.635 0.321 0.478
XGBoost (Iterative-Top 5) 0.634 0.301 0.467
F.8 Time Benchmark
We benchmark the time elapsed to run training and performing inference. All experiments are run on the same machine with AMD EPYC 9354 32-Core Processor and NVIDIA RTX 6000 Ada GPU with 48GB memory. We show the detailed runtime for train and test split, and calibration on SummEval dataset in Table 20. Specifically, for G-Eval (GPT4), it costs around US\$82 to run SummEval benchmark with a max token of 5, temperature of 2, top p of 1, a frequency penalty of 0, a presence penalty of 0, and completion choices of 20.
To compare the proper time evaluation on test split with MetaMetrics-SUM, it is important to account for all the components of MetaMetrics-SUM’s evaluation process. Specifically, for individual metrics, we can consider the inference time on the test split. On the other hand, for MetaMetrics-SUM, we need to consider the total time, consisting of inference time on both the train and test splits, and then the calibration time using the train set. For example, G-Eval (GPT4) takes 7,173.50 seconds to evaluate on the test split, while XGBoost (all metrics including LLM-based metrics) takes 3,976.47 $+$ 9,292.83 $+$ 647.08 $=$ 13,916.38 seconds to evaluate on the test split.
Certain metric groups, such as BLEURT, SummaQA, and BERTScore, have multiple variations, all of which can be computed in a single evaluation pass. Therefore, the time measurement for MetaMetrics-SUM reflects the total duration of the pass, rather than the summation of time of all individual rows.
An important observation is that the bottleneck for MetaMetrics-SUM lies in the inference time for both the train and test splits, which dominates the total runtime compared to the calibration time. This is particularly evident when MetaMetrics-SUM includes metrics like G-Eval. The calibration time is relatively minor, with the longest calibration being for GP (all metrics with LLM-based metrics), which takes 647.08 seconds. However, if the metric outputs are already available, the inference step can be skipped.
Currently, the reported time for processing the train and test splits for MetaMetrics-SUM assumes a sequential evaluation of metrics. However, this sequential approach can be optimized by running metrics in parallel, as their evaluations are completely independent of each other. In other words, these computations are embarrassingly parallelizable. With this optimization, the total runtime for MetaMetrics-SUM would effectively be the time taken by the slowest metric, plus the calibration time. For example, for GP (all metrics except LLM-based metrics), the longest metric that takes to be evaluated is SummaQA, which takes 790.11 seconds. Therefore, the total running time when running MetaMetrics-NUM in a completely parallel manner will be 790.11 $+$ 402.51 $=$ 1,192.62 seconds.
Table 20: Runtime for inference time on train and test split and the calibration time for MetaMetrics. Rows with exact same run time is due to the groups of metrics (e.g., BLEURT (Max) and BLEURT (Mean)) being calculated in a single pass. G-Eval (GPT4) costs around US\$82 to run SummEval benchmark with a max token of 5, temperature of 2, top p of 1, a frequency penalty of 0, a presence penalty of 0, and completion choices of 20. The running time reported for MetaMetrics-SUM for train and test split is in a sequential manner. In other words, running each metric in an embarrassingly parallel way can reduce the overall inference time. To compare the proper time evaluation on test split with MetaMetrics-SUM, it is important to account for all the components of MetaMetrics-SUM’s evaluation process. Specifically, for individual metrics, we can consider the inference time on the test split. On the other hand, for MetaMetrics-SUM, we need to consider the inference time on both the train and test splits, and then the calibration time using the train set. For example, G-Eval (GPT4) takes 7,173.50 seconds to evaluate on the test split, while XGBoost (all metrics including LLM-based metrics) takes 3,976.47 $+$ 9,292.83 $+$ 647.08 $=$ 13,916.38 seconds to evaluate on the test split.
Metric Inference Time (in sec) Calibration Time (in sec)
Train Split Test Split
BLEU 0.72 1.73 N/A
chrF 3.37 8.67 N/A
METEOR 5.74 13.57 N/A
ROUGE-1 57.19 162.77 N/A
ROUGE-3 68.24 151.90 N/A
ROUGE-4 70.43 149.65 N/A
ROUGE-l 66.53 155.15 N/A
ROUGE-WE1 56.35 100.19 N/A
BLEURT (Max) 139.87 336.22 N/A
BLEURT (Mean) 139.87 336.22 N/A
BERTScore (f1) 1.74 3.99 N/A
BERTScore (Recall) 1.74 3.99 N/A
SummaQA (f1) 322.07 790.11 N/A
SummaQA (Prob) 322.07 790.11 N/A
LLM-based Metrics
BARTScore (Max) 22.39 50.83 N/A
BARTScore (Mean) 22.39 50.83 N/A
UniEval 80.37 194.55 N/A
G-Eval (GPT4) 3,081.46 7,173.50 N/A
Ensemble Baselines
Uniform 3,976.47 9,292.83 N/A
Weighted Avg 3,976.47 9,292.83 N/A
MetaMetrics-Sum
   GP 792.25 1,873.95 402.51
   XGBoost 792.25 1,873.95 276.52
w/ LLM-based Metrics
   GP 3,976.47 9,292.83 647.08
   GP (Top 2) 3,161.83 7,368.05 187.73
   XGBoost 3,976.47 9,292.83 337.88
   XGBoost (Top 2) 3,161.83 7,368.05 223.69
Appendix G Weights and Feature Importance
G.1 Question Answering